COST OF EQUITY WITH AND WITHOUT FLOTATION

Jarett & Sons's common stock currently trades at $39.00 a share. It is expected to pay an annual dividend of $3.00 a share at the end of the year (D1 = $3.00), and the constant growth rate is 5% a year.

  1. What is the company's cost of common equity if all of its equity comes from retained earnings? Round your answer to two decimal places. Do not round your intermediate calculations.
    %

  2. If the company issued new stock, it would incur a 8% flotation cost. What would be the cost of equity from new stock? Round your answer to two decimal places. Do not round your intermediate calculations.
    %
Answer #1

a.Cost of equity=(D1/Current price)+Growth rate

=(3/39)+0.05

=12.69%(Approx).

b.Cost of equity=(D1/Current price(1-Flotation cost)+Growth rate

=(3/39(1-0.08))+0.05

=(3/35.88)+0.05

=13.36%(Approx).